# Integrating Dynamical Systems Modeling with Spatiotemporal scRNA-seq Data Analysis

### v1

@@ Abstract
Single-cell RNA sequencing, Spatiotemporal dynamics, Computational modeling, Cellular trajectories.
Contents
@@ 1 Introduction
Understanding the dynamic change of biological systems has played a central role in life sciences, with important applications in developmental biology, disease modeling, and medicine (Lei, 2023; Hong and Xing, 2024; Xing, 2022; Schiebinger, 2021; Heitz et al., 2024). One key framework for understanding these dynamic processes is Waddington’s developmental landscape (Waddington, 2014; Moris et al., 2016; MacLean et al., 2018), which illustrates how cells navigate various potential fates as they differentiate during development. However, how to construct such developmental landscapes or understand the cellular dynamics within the biological systems, presents a significant challenge. To fully understand these complex cellular transitions, a deep understanding of gene expression at the single-cell level is essential. Advancements in high-throughput sequencing technologies have enabled unprecedented resolutions into the molecular signatures of individual cells, with single-cell RNA sequencing (scRNA-seq) emerging as a revolutionary tool. scRNA-seq allows for the dissection of cellular heterogeneity and the identification of transcriptional programs underlying complex biological processes, offering a snapshot of gene expression in single cells at a given moment. Despite its powerful capabilities, traditional scRNA-seq provides only a static picture of gene expression across individual cells, missing the temporal information for understanding how cells transition through different states.
In recent years, the development of temporally resolved scRNA-seq technologies has begun to gain increasing attention, enabling the capture of gene expression profiles across multiple time points. Another breakthrough in transcriptomics is spatial transcriptomics (ST), which integrates spatial context into gene expression data by mapping RNA profiles within tissue architectures. When combined with temporal resolution, this approach leads to temporally resolved spatial transcriptomics (temporal-ST), which provides an enhanced tool for studying the spatiotemporal dynamics of single cells.
@@
To address the challenges numerous computational frameworks have been developed. For single-cell transcriptomics data, several methods have been proposed to approximate cellular trajectories and dissect dynamic cellular states. Pseudotime inference methods (Qiu et al., 2017; Cao et al., 2019; Street et al., 2018), for instance, arrange snapshot data along an inferred developmental axis, offering a continuous perspective of cell-state transitions over time. In addition, RNA velocity analysis (La Manno et al., 2018; Bergen et al., 2021; Wang et al., 2024a; Liu et al., 2024; Bergen et al., 2020) has emerged as a powerful tool for understanding cellular dynamics by leveraging splicing kinetics to infer the direction of future gene expression changes. Recently, with the development of temporally resolved sequencing technology, there has been a growing interest in dissecting single-cell dynamics from multiple snapshot data. Simultaneously, the development of generative modeling techniques, such as diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015; Song et al., 2021; Ren et al., 2025), optimal transport theory (Bunne et al., 2024; Schiebinger et al., 2019; Klein et al., 2025), flow-based model (Lipman et al., 2023; Tong et al., 2024a), and the Schrödinger bridge problem (Gentil et al., 2017; Zhang et al., 2025) have emerged as key mathematical frameworks for modeling distribution transitions in dynamic biological systems. More recently, the rapid development of spatial transcriptomics has opened exciting new avenues for integrating spatial and temporal data. Extending these computational methods to the ST data and capturing spatiotemporal cellular transitions also inspired many recent kinds of research (Peng et al., 2024).
In recent years, many reviews have provided comprehensive summaries of the methods and advancements in the study of single-cell dynamics. For example, (Wang et al., 2021b) reviewed various pseudotime inference methods. (Saelens et al., 2019) conducted a comprehensive benchmarking study on pseudotime inference methods. (Wang et al., 2021b; Li et al., 2020; Bergen et al., 2021), reviewed RNA velocity methods in single-cell transcriptomics. (Bergen et al., 2021) also discussed the limitations and potential extensions of RNA velocity. (Bunne et al., 2024; Heitz et al., 2024; Schiebinger, 2021; Jiang Qi, 2025) provided an in-depth analysis of the application of optimal transport theory in single-cell or spatial omics data. Additionally, (Xing, 2022; Lei, 2023) examined various perspectives on cellular dynamics, exploring how the reconstruction of cell states and energy landscapes can contribute to our understanding of cellular behavior and development. The current review takes a distinct perspective by systematically discussing modeling strategies for different types of data from a dynamical modeling perspective, aiming to unify and expand upon the current methodologies in the field.
@@
Figure 1: Overview of the Data and Models. (a) Discrete and Continuous Model: The discrete model constructs Markov chains between cells with dynamics encoded in a transition matrix, while the continuous model describes single-cell motion via stochastic differential equations (SDEs) and cell population dynamics through a corresponding partial differential equation (PDE). (b) Datasets: Snapshot data is an $n\times g$ matrix $\mathbf{X}$ ($n$: cell count, $g$: gene count); Temporal data provides gene expression matrices $\mathbf{X}_{i}$ at time points $i\in\{0,\ldots,T-1\}$; Spatial data additionally records coordinates for each cell in $\mathbf{X}_{i}$.
@@ 2 Overview of the Data and Models
In this section, we provide preliminary background on the structure of the scRNA-seq data as well as the mathematical models to describe dynamic cellular processes.
@@ 2.1.1 Snapshot scRNA-seq Data
In snapshot RNA sequencing (RNA-seq) data, gene expression is measured across multiple cells at a single time point. The gene expression matrix is represented as $\mathbf{X}\in\mathbb{R}^{n\times d}$, where $\mathbf{X}$ denotes the count matrix of gene expression, $n$ is the number of cells or spots, and $d$ is the number of genes measured. Each entry $X_{ij}$ in $\mathbf{X}$ represents the expression level of gene $j$ in cell $i$, typically measured as the number of mRNA molecules (transcripts) for that gene in the corresponding cell. Additionally, the total RNA-seq data can further be separated into counts for spliced and unspliced transcripts, useful in certain analyses such as the RNA velocity model described below. The spliced and unspliced counts are represented as $\mathbf{U}\in\mathbb{R}^{n\times d}$ and $\mathbf{S}\in\mathbb{R}^{n\times d}$, respectively, denoting the matrices of unspliced and spliced counts for the $n$ cells or spots. Over time, unspliced RNA ($\mathbf{u}$) can undergo splicing process to become spliced RNA ($\mathbf{s}$).
@@ 2.1.2 Temporally and Spatially Resolved scRNA-seq
Recently, a growing number of temporally resolved scRNA-seq datasets have been generated, where single-cell measurements are performed at multiple time points during a dynamic process. Such datasets could offer deeper insights into how cell populations evolve over time (Ding et al., 2022; Schiebinger, 2021; Bunne et al., 2024).
For temporally resolved scRNA-seq dataset, at each fixed time point $i\in\{0,\dots,T-1\}$, the gene expression matrix is represented as $\mathbf{X}_{i}\in\mathbb{R}^{n_{i}\times d}$, where $\mathbf{X}_{i}$ denotes the matrix of gene expression data, $n_{i}$ is the number of cells at time $i$, and $d$ is the number of genes. Notably, the gene expression data across time points are unpaired and can be assumed to be sampled from a distribution from a certain time point.
The development of spatial transcriptomics (ST) technology allows gene expression to be captured alongside spatial coordinates (Ståhl et al., 2016; Rodriques et al., 2019; Stickels et al., 2021; Chen et al., 2022a; Oliveira et al., 2024; Moffitt et al., 2018; Eng et al., 2019; Wang et al., 2018). ST methods are broadly divided into image-based and sequencing-based approaches. Image-based techniques (Moffitt et al., 2018; Eng et al., 2019; Wang et al., 2018) detect hundreds to thousands of genes with cellular or sub-cellular resolution, while sequencing-based methods (Rodriques et al., 2019; Stickels et al., 2021; Chen et al., 2022a; Oliveira et al., 2024) allow whole-transcriptome analysis but are usually limited to spot-level resolution. Advances like Stereo-seq (Chen et al., 2022a) and 10x Visium HD (Oliveira et al., 2024) have significantly improved spatial resolution to single-cell or even subcellular precision.
Similarly to temporally resolved scRNA-seq data, ST time series data could be represented as $(\mathbf{Z}_{(0:K)},\mathbf{X}_{(0:K)})$ at $t_{0},t_{1}\dots t_{K}$ totaling $K$ time points, and the number of cells in each observation is $n_{0},n_{1}\dots n_{K}$. In addition to the gene expression matrices $\mathbf{X}_{i}\in\mathbb{R}^{n_{i}\times d}$, the associated spatial coordinate matrices $\mathbf{Z}_{i}\in\mathbb{R}^{n_{i}\times 2}$ or $\mathbb{R}^{n_{i}\times 3}$ represent the spatial coordinates (2D or 3D) of each sequenced cell (or spot) respectively.
@@ 2.2.1 Discrete Dynamics: Markov Chain Model
Random walk or Markov chain models are simple yet powerful tools for studying stochastic dynamic processes, particularly in the context of complex systems such as gene expression dynamics and cell trajectories. In these models, a system evolves over time as a series of transitions between discrete states, where each state corresponds to a possible configuration or position in the system (such as a specific gene expression profile or a cell’s position in a developmental trajectory). The transitions between states are governed by transition probabilities, which can be represented in the form of a transition matrix $\mathbf{P}$. The transition matrix is defined as $P_{ij}=\frac{W_{ij}}{W_{i}},$ where $P_{ij}$ represents the probability of transitioning from cell state $i$ to state $j$, and $W_{ij}$ is the weight (or similarity) between cells $i$ and $j$ in a weighted graph, and $W_{i}=\sum_{k}W_{ik}$ is the degree of cell $i$. The weights $W_{ij}$ typically reflect some measure of similarity or distance between the corresponding gene expression profiles of the cells, or induced from other quantities such as RNA velocity or optimal transport plan.
@@
A more realistic assumption in biology is that the state-transition graph can be directional (for example, induced by RNA velocity or optimal transport discussed below), with the cells ultimately reaching terminal states such as fully differentiated or mature cell types. In this setup, recurrent states represent the final, stable cell types or fates that the system eventually reaches. Once a cell enters one of these recurrent states, it remains there, similar to how a fully differentiated cell does not revert back to an undifferentiated or less specialized state. On the other hand, transient states correspond to intermediate stages of cellular development, such as precursor or progenitor cells that are still undergoing differentiation or division. These cells are in transition, with the potential to eventually reach one of the recurrent, stable cell types. The transition matrix governing this system can be partitioned into blocks that reflect these different types of cell states. Specifically, the matrix $\mathbf{P}$ can be written as the canonical form
$P=\begin{bmatrix}{\mathbf{\tilde{P}}}&\mathbf{0}\\ \mathbf{S}&\mathbf{Q}\end{bmatrix}.$ (2.1)
Here, $\tilde{\mathbf{P}}$ corresponds to the transitions between recurrent (terminal) states, where once a cell reaches these states, it remains there in absorbing states. $\mathbf{Q}$ represents transitions between transient states (cells that are still in intermediate stages of differentiation or cell cycle). $\mathbf{S}$ denotes the transitions from transient to recurrent states, representing cells’ eventual differentiation or maturation into their terminal stable types. Since recurrent states are absorbing, the upper right block of the matrix is zero, indicating no transitions from recurrent to transient states.
@@ 2.2.2 Continuous Dynamics: From Trajectories to Population Dynamics
Trajectory Dynamics: Stochastic Differential Equations (SDEs)
@@
To model the evolution of individual cell trajectories, we consider that cellular dynamics can be governed by a stochastic differential equation (SDE). This accounts for both deterministic factors, such as gene expression regulation, and stochastic factors, like noise from cellular environments or molecular fluctuations. The SDE for the state $\mathbf{x}_{t}$ of a single cell at time $t$ is given by
$\mathrm{d}\mathbf{x}_{t}=\mathbf{b}(\mathbf{x}_{t},t)\,\mathrm{d}t+\boldsymbol{\sigma}(\mathbf{x}_{t},t)\,\mathrm{d}\mathbf{w}_{t},$ (2.2)
@@
where $\mathbf{x}_{t}\in\mathbb{R}^{d}$ represents the state of the cell (e.g., gene expression profile) at time $t$, and $\mathbf{w}_{t}\in\mathbb{R}^{d}$ is the standard $d$-dimensional Brownian motion. The term $\mathbf{b}(\mathbf{x},t)$ represents the drift vector, which defines the deterministic flow of the system, while $\boldsymbol{\sigma}(\mathbf{x},t)\in\mathbb{R}^{d\times d}$ represents the diffusion coefficient matrix, which describes the random fluctuations in the system.
@@
Specifically, when the diffusion coefficient $\boldsymbol{\sigma}(\mathbf{x},t)$ is zero, the system reduces to an ordinary differential equation (ODE), which describes the deterministic evolution of the cell state without random fluctuations. In this case, the evolution of the cell is entirely governed by the drift term $\mathbf{b}(\mathbf{x},t)$, and the system follows a deterministic trajectory. A useful concept in understanding the long-term behavior of the system is that of an attractor. In the context of the ODE, an attractor corresponds to a stable fixed point of the system, where the rate of change $\mathbf{b}(\mathbf{x},t)$ of the cell state $\mathbf{x}$ becomes zero. In cellular dynamics, such attractors can represent stable gene expression profiles, such as differentiated or quiescent states, where the cell remains in a stable state over time.
@@
Furthermore, since $\mathbf{b}(\mathbf{x},t)$ is time-dependent, the system may exhibit bifurcations, where the qualitative characteristics of attractors could change with respect to time $t$. In cellular contexts, bifurcations are important for understanding processes like cell fate decisions, where a small change in the environment or internal signaling can push the cell toward a new distinct state (e.g., differentiation into a different cell type).
Population Dynamics: Partial Differential Equations (PDEs)
@@
To capture the evolution of the entire population of cells, we consider the density of cells with respect to their state $\mathbf{x}$, represented by the probability density function $p(\mathbf{x},t)$. The population distribution evolves according to a partial differential equation (PDE), which incorporates both deterministic and stochastic dynamics at the population level. As a simple model, the evolution of $p(\mathbf{x},t)$ can be described by
$\partial_{t}p(\mathbf{x},t)=-\nabla_{\mathbf{x}}\cdot\left(p(\mathbf{x},t)\,\mathbf{b}(\mathbf{x},t)\right)+\frac{1}{2}\nabla_{\mathbf{x}}^{2}:\left(\mathbf{a}(\mathbf{x},t)\,p(\mathbf{x},t)\right)+g(\mathbf{x},t)\,p(\mathbf{x},t),$ (2.3)
@@
where $\nabla_{\mathbf{x}}^{2}:\left(\mathbf{a}(\mathbf{x},t)\,p(\mathbf{x},t)\right)=\sum_{ij}\partial_{ij}\left(a_{ij}(\mathbf{x},t)p(\mathbf{x},t)\right)$, and $\mathbf{a}(\mathbf{x},t)=\boldsymbol{\sigma}(\mathbf{x},t)\boldsymbol{\sigma}^{T}(\mathbf{x},t)$ represents the diffusion matrix at the population level.
@@
The terms on the right-hand side of the equation represent the key dynamics driving the population evolution. The drift term $\nabla_{\mathbf{x}}\cdot\left(p(\mathbf{x},t)\,\mathbf{b}(\mathbf{x},t)\right)$ quantifies the deterministic flow of the population, describing how cells move through different states based on the drift vector $\mathbf{b}(\mathbf{x},t)$. The diffusion term $\frac{1}{2}\nabla_{\mathbf{x}}^{2}:\left(\mathbf{a}(\mathbf{x},t)\,p(\mathbf{x},t)\right)$ models the spread of the population due to random fluctuations, where $\mathbf{a}(\mathbf{x},t)=\boldsymbol{\sigma}(\mathbf{x},t)\boldsymbol{\sigma}^{T}(\mathbf{x},t)$ represents the diffusion matrix, capturing the effects of stochasticity. Finally, the growth term $g(\mathbf{x},t)\,p(\mathbf{x},t)$ governs the birth and death rates of cells, modeling cell proliferation and mortality, thus controlling the population’s size and dynamics over time. When $g(\mathbf{x},t)=0$, the PDE reduces to the Fokker-Planck equation associated with the SDE in the Itô integral sense, describing the evolution of the probability density $p(\mathbf{x},t)$ for the stochastic process defined by the SDE.

### v2

# single-cell RNA sequencing; spatiotemporal dynamics; computational modeling; cellular trajectories

@@ Abstract
single-cell RNA sequencing; spatiotemporal dynamics; computational modeling; cellular trajectories
[…]
@@ 1 Introduction
Understanding the dynamic change of biological systems has played a central role in life sciences, with important applications in developmental biology, disease modeling, and medicine Lei (2023); Hong and Xing (2024); Xing (2022); Schiebinger (2021); Heitz et al. (2024). One key framework for understanding these dynamic processes is Waddington’s developmental landscape Waddington (2014); Moris et al. (2016); MacLean et al. (2018), which illustrates how cells navigate various potential fates as they differentiate during development. However, how to construct such developmental landscapes or understand the cellular dynamics within the biological systems, presents a significant challenge. To fully understand these complex cellular transitions, a deep understanding of gene expression at the single-cell level is essential. Advancements in high-throughput sequencing technologies have enabled unprecedented resolutions into the molecular signatures of individual cells, with single-cell RNA sequencing (scRNA-seq) emerging as a revolutionary tool Ziegenhain et al. (2017); Tang et al. (2009); Stark et al. (2019). scRNA-seq allows for the dissection of cellular heterogeneity and the identification of transcriptional programs underlying complex biological processes, offering a snapshot of gene expression in single cells at a given moment. Despite its powerful capabilities, traditional scRNA-seq provides only a static picture of gene expression across individual cells, missing the temporal information for understanding how cells transition through different states.
In recent years, the development of temporally resolved scRNA-seq technologies has begun to gain increasing attention, enabling the capture of gene expression profiles across multiple time points Ding et al. (2022); Schiebinger (2021); Bunne et al. (2024). Another breakthrough in transcriptomics is spatial transcriptomics (ST), which integrates spatial context into gene expression data by mapping RNA profiles within tissue architectures Ståhl et al. (2016); Rodriques et al. (2019); Stickels et al. (2021); Chen et al. (2022); Oliveira et al. (2024); Moffitt et al. (2018); Eng et al. (2019); Wang et al. (2018). When combined with temporal resolution, this approach leads to temporally resolved spatial transcriptomics (temporal-ST), which provides an enhanced tool for studying the spatiotemporal dynamics of single cells Liu et al. (2024).
@@
To address the challenges, numerous computational frameworks have been developed. For single-cell transcriptomics data, several methods have been proposed to approximate cellular trajectories and dissect dynamic cellular states. Pseudotime inference methods Qiu et al. (2017); Cao et al. (2019); Street et al. (2018), for instance, arrange snapshot data along an inferred developmental axis, offering a continuous perspective of cell-state transitions over time. In addition, RNA velocity analysis La Manno et al. (2018); Bergen et al. (2021); Wang et al. (2024); Liu et al. (2024); Bergen et al. (2020) has emerged as a powerful tool for understanding cellular dynamics by leveraging splicing kinetics to infer the direction of future gene expression changes. Recently, with the development of temporally resolved sequencing technology, there has been a growing interest in dissecting single-cell dynamics from multiple snapshot data. Simultaneously, the development of generative modeling techniques, such as diffusion models Ho et al. (2020); Sohl-Dickstein et al. (2015); Song et al. (2021); Ren et al. (2025), optimal transport theory Bunne et al. (2024); Schiebinger et al. (2019); Klein et al. (2025), flow-based model Lipman et al. (2023); Tong et al. (2024), and the Schrödinger bridge problem Gentil et al. (2017); Zhang et al. (2025) have emerged as key mathematical frameworks for modeling distribution transitions in dynamic biological systems. More recently, the rapid development of spatial transcriptomics has opened exciting new avenues for integrating spatial and temporal data. Extending these computational methods to the ST data and capturing spatiotemporal cellular transitions also has inspired many recent kinds of research Peng et al. (2024).
Recently, many reviews have provided comprehensive summaries of the methods and advancements in the study of single-cell dynamics. For example, ref. Wang et al. (2021) reviewed various pseudotime inference methods. Ref. Saelens et al. (2019) conducted a comprehensive benchmarking study on pseudotime inference methods.  Refs. Wang et al. (2021); Li et al. (2020); Bergen et al. (2021) reviewed RNA velocity methods in single-cell transcriptomics. Bergen et al. (2021) also discussed the limitations and potential extensions of RNA velocity. Refs. Bunne et al. (2024); Heitz et al. (2024); Schiebinger (2021); Jiang Qi (2025) provided an in-depth analysis of the application of optimal transport theory in single-cell or spatial omics data. Additionally, refs. Xing (2022); Lei (2023) examined various perspectives on cellular dynamics, exploring how the reconstruction of cell states and energy landscapes can contribute to our understanding of cellular behavior and development. The current review takes a distinct perspective by systematically discussing modeling strategies for different types of data from a dynamical modeling perspective, aiming to unify and expand upon the current methodologies in the field.
@@ 2 Overview of the Data and Models
In this section, we provide preliminary background on the structure of the scRNA-seq data as well as the mathematical models to describe dynamic cellular processes. An overview of the data and models is provided in Figure 1.
{adjustwidth}
[…]
Figure 1: Overview of the data and models. (a) Discrete and Continuous Model: The discrete model constructs Markov chains between cells with dynamics encoded in a transition matrix, while the continuous model describes single-cell motion via stochastic differential equations (SDEs) and cell population dynamics through a corresponding partial differential equation (PDE). (b) Datasets: Snapshot data is an $n\times g$ matrix $\boldsymbol{X}$ ($n$: cell count, $g$: gene count); temporal data provides gene expression matrices $\boldsymbol{X}_{i}$ at time points $i\in\{0,\ldots,T-1\}$; spatial data additionally records coordinates for each cell in $\boldsymbol{X}_{i}$.
@@ 2.1.1 Snapshot scRNA-Seq Data
In snapshot RNA sequencing (RNA-seq) data, gene expression is measured across multiple cells at a single time point. The gene expression matrix is represented as $\boldsymbol{X}\in\mathbb{R}^{n\times d}$, where $\boldsymbol{X}$ denotes the count matrix of gene expression, $n$ is the number of cells or spots, and $d$ is the number of genes measured. Each entry $X_{ij}$ in $\boldsymbol{X}$ represents the expression level of gene $j$ in cell $i$, typically measured as the number of mRNA molecules (transcripts) for that gene in the corresponding cell. Additionally, the total RNA-seq data can further be separated into counts for spliced and unspliced transcripts, useful in certain analyses such as the RNA velocity model described below. The spliced and unspliced counts are represented as $\boldsymbol{U}\in\mathbb{R}^{n\times d}$ and $\boldsymbol{S}\in\mathbb{R}^{n\times d}$, respectively, denoting the matrices of unspliced and spliced counts for the $n$ cells or spots. Over time, unspliced RNA ($\boldsymbol{u}$) can undergo splicing process to become spliced RNA ($\boldsymbol{s}$).
@@ 2.1.2 Temporally and Spatially Resolved scRNA-Seq
Recently, a growing number of temporally resolved scRNA-seq datasets have been generated, where single-cell measurements are performed at multiple time points during a dynamic process. Such datasets could offer deeper insights into how cell populations evolve over time Ding et al. (2022); Schiebinger (2021); Bunne et al. (2024).
For temporally resolved scRNA-seq dataset, at each fixed time point $i\in\{0,\dots,T-1\}$, the gene expression matrix is represented as $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{i}\times d}$, where $\boldsymbol{X}_{i}$ denotes the matrix of gene expression data, $n_{i}$ is the number of cells at time $i$, and $d$ is the number of genes. Notably, the gene expression data across time points are unpaired and can be assumed to be sampled from a distribution from a certain time point.
The development of spatial transcriptomics (ST) technology allows gene expression to be captured alongside spatial coordinates  Ståhl et al. (2016); Rodriques et al. (2019); Stickels et al. (2021); Chen et al. (2022); Oliveira et al. (2024); Moffitt et al. (2018); Eng et al. (2019); Wang et al. (2018). ST methods are broadly divided into image-based and sequencing-based approaches. Image-based techniques  Moffitt et al. (2018); Eng et al. (2019); Wang et al. (2018) detect hundreds to thousands of genes with cellular or sub-cellular resolution, while sequencing-based methods  Rodriques et al. (2019); Stickels et al. (2021); Chen et al. (2022); Oliveira et al. (2024) allow whole-transcriptome analysis but are usually limited to spot-level resolution. Advances like Stereo-seq Chen et al. (2022) and 10x Visium HD Oliveira et al. (2024) have significantly improved spatial resolution to single-cell or even subcellular precision.
Similarly to temporally resolved scRNA-seq data, ST time series data could be represented as $(\boldsymbol{Z}_{(0:K)},\boldsymbol{X}_{(0:K)})$ at $t_{0},t_{1}\dots t_{K}$ totaling $K$ time points, and the number of cells in each observation is $n_{0},n_{1}\dots n_{K}$. In addition to the gene expression matrices $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{i}\times d}$, the associated spatial coordinate matrices $\boldsymbol{Z}_{i}\in\mathbb{R}^{n_{i}\times 2}$ or $\mathbb{R}^{n_{i}\times 3}$ represent the spatial coordinates (2D or 3D) of each sequenced cell (or spot), respectively.
@@ 2.2.1 Discrete Dynamics: Markov Chain Model
Random walk or Markov chain models are simple yet powerful tools for studying stochastic dynamic processes, particularly in the context of complex systems such as gene expression dynamics and cell trajectories. In these models, a system evolves over time as a series of transitions between discrete states, where each state corresponds to a possible configuration or position in the system (such as a specific gene expression profile or a cell’s position in a developmental trajectory). The transitions between states are governed by transition probabilities, which can be represented in the form of a transition matrix $\boldsymbol{P}$. The transition matrix is defined as $P_{ij}=\frac{W_{ij}}{W_{i}},$ where $P_{ij}$ represents the probability of transitioning from cell state $i$ to state $j$, and $W_{ij}$ is the weight (or similarity) between cells $i$ and $j$ in a weighted graph, and $W_{i}=\sum_{k}W_{ik}$ is the degree of cell $i$. The weights $W_{ij}$ typically reflect some measure of similarity or distance between the corresponding gene expression profiles of the cells, or induced from other quantities such as RNA velocity or optimal transport plan.
@@
A more realistic assumption in biology is that the state-transition graph can be directional (for example, induced by RNA velocity or optimal transport discussed below), with the cells ultimately reaching terminal states such as fully differentiated or mature cell types. In this setup, recurrent states represent the final, stable cell types or fates that the system eventually reaches. Once a cell enters one of these recurrent states, it remains there, similar to how a fully differentiated cell does not revert back to an undifferentiated or less specialized state. On the other hand, transient states correspond to intermediate stages of cellular development, such as precursor or progenitor cells that are still undergoing differentiation or division. These cells are in transition, with the potential to eventually reach one of the recurrent, stable cell types. The transition matrix governing this system can be partitioned into blocks that reflect these different types of cell states. Specifically, the matrix $\boldsymbol{P}$ can be written as the canonical form
$\boldsymbol{P}=\begin{bmatrix}{\boldsymbol{\tilde{P}}}&\boldsymbol{0}\\ \boldsymbol{S}&\boldsymbol{Q}\end{bmatrix}.$ (1)
Here, $\tilde{\boldsymbol{P}}$ corresponds to the transitions between recurrent (terminal) states, where once a cell reaches these states, it remains there in absorbing states. $\boldsymbol{Q}$ represents transitions between transient states (cells that are still in intermediate stages of differentiation or cell cycle). $\boldsymbol{S}$ denotes the transitions from transient to recurrent states, representing cells’ eventual differentiation or maturation into their terminal stable types. Since recurrent states are absorbing, the upper right block of the matrix is zero, indicating no transitions from recurrent to transient states.
@@ 2.2.2 Continuous Dynamics: From Trajectories to Population Dynamics
Trajectory Dynamics: Stochastic Differential Equations (SDEs) To model the evolution of individual cell trajectories, we consider that cellular dynamics can be governed by a stochastic differential equation (SDE). This accounts for both deterministic factors, such as gene expression regulation, and stochastic factors, like noise from cellular environments or molecular fluctuations. The SDE for the state $\boldsymbol{x}_{t}$ of a single cell at time $t$ is given by
$\mathrm{d}\boldsymbol{x}_{t}=\boldsymbol{b}(\boldsymbol{x}_{t},t)\,\mathrm{d}t+\boldsymbol{\sigma}(\boldsymbol{x}_{t},t)\,\mathrm{d}\boldsymbol{w}_{t},$ (2)
@@
where $\boldsymbol{x}_{t}\in\mathbb{R}^{d}$ represents the state of the cell (e.g., gene expression profile) at time $t$, and $\boldsymbol{w}_{t}\in\mathbb{R}^{d}$ is the standard $d$-dimensional Brownian motion. The term $\boldsymbol{b}(\boldsymbol{x},t)$ represents the drift vector, which defines the deterministic flow of the system, while $\boldsymbol{\sigma}(\boldsymbol{x},t)\in\mathbb{R}^{d\times d}$ represents the diffusion coefficient matrix, which describes the random fluctuations in the system.
@@
Specifically, when the diffusion coefficient $\boldsymbol{\sigma}(\boldsymbol{x},t)$ is zero, the system reduces to an ordinary differential equation (ODE), which describes the deterministic evolution of the cell state without random fluctuations. In this case, the evolution of the cell is entirely governed by the drift term $\boldsymbol{b}(\boldsymbol{x},t)$, and the system follows a deterministic trajectory. A useful concept in understanding the long-term behavior of the system is that of an attractor.  In the context of the ODE, an attractor corresponds to a stable fixed point of the system, where the rate of change $\boldsymbol{b}(\boldsymbol{x},t)$ of the cell state $\boldsymbol{x}$ becomes zero. In cellular dynamics, such attractors can represent stable gene expression profiles, such as differentiated or quiescent states, where the cell remains in a stable state over time.
@@
Furthermore, since $\boldsymbol{b}(\boldsymbol{x},t)$ is time-dependent, the system may exhibit bifurcations, where the qualitative characteristics of attractors could change with respect to time $t$. In cellular contexts, bifurcations are important for understanding processes like cell fate decisions, where a small change in the environment or internal signaling can push the cell toward a new distinct state (e.g., differentiation into a different cell type).
@@
Population Dynamics: Partial Differential Equations (PDEs) To capture the evolution of the entire population of cells, we consider the density of cells with respect to their state $\boldsymbol{x}$, represented by the probability density function $p(\boldsymbol{x},t)$. The population distribution evolves according to a partial differential equation (PDE), which incorporates both deterministic and stochastic dynamics at the population level. As a simple model, the evolution of $p(\boldsymbol{x},t)$ can be described by
$\partial_{t}p(\boldsymbol{x},t)=-\nabla_{\boldsymbol{x}}\cdot\left(p(\boldsymbol{x},t)\,\boldsymbol{b}(\boldsymbol{x},t)\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t)\,p(\boldsymbol{x},t)\right)+g(\boldsymbol{x},t)\,p(\boldsymbol{x},t),$ (3)
@@
where $\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t)\,p(\boldsymbol{x},t)\right)=\sum_{ij}\partial_{ij}\left(a_{ij}(\boldsymbol{x},t)p(\boldsymbol{x},t)\right)$, and $\boldsymbol{a}(\boldsymbol{x},t)=\boldsymbol{\sigma}(\boldsymbol{x},t)\boldsymbol{\sigma}^{T}(\boldsymbol{x},t)$ represents the diffusion matrix at the population level.
@@
The terms on the right-hand side of the equation represent the key dynamics driving the population evolution. The drift term $\nabla_{\boldsymbol{x}}\cdot\left(p(\boldsymbol{x},t)\,\boldsymbol{b}(\boldsymbol{x},t)\right)$ quantifies the deterministic flow of the population, describing how cells move through different states based on the drift vector $\boldsymbol{b}(\boldsymbol{x},t)$. The diffusion term $\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t)\,p(\boldsymbol{x},t)\right)$ models the spread of the population due to random fluctuations, where $\boldsymbol{a}(\boldsymbol{x},t)=\boldsymbol{\sigma}(\boldsymbol{x},t)\boldsymbol{\sigma}^{T}(\boldsymbol{x},t)$ represents the diffusion matrix, capturing the effects of stochasticity. Finally, the  growth term $g(\boldsymbol{x},t)\,p(\boldsymbol{x},t)$ governs the birth and death rates of cells, modeling cell proliferation and mortality, thus controlling the population’s size and dynamics over time. When $g(\boldsymbol{x},t)=0$, the PDE reduces to the Fokker–Planck equation associated with the SDE in the Itô integral sense, describing the evolution of the probability density $p(\boldsymbol{x},t)$ for the stochastic process defined by the SDE.